10. A high-tech company has invested 15% of its assets in financial stocks. Suppose that we select 5 stocks at random (i)What is the probability that at least 3 of the stocks are financial stocks? (ii)What is the probability that none of the stocks are financial stocks?
Answer #1

X ~ Binomial (n,p)

Where n = 5 , p = 0.15

Binomial probability distribution is

P(X) = nCx px (1-p)n-x

a)

P( X >= 3) = P( x = 3) + P( x = 4) + P( X = 5)

= 5C3 0.153 0.852 + 5C4 0.154 0.85 + 5C5 0.155 0.850

= 0.0266

b)

P( X = 0) = 5C0 0.150 0.855

= 0.4437